# Decoherence-free interaction and maximally entangled state generation in double-giant-atom semi-infinite waveguide systems

### v1

@@ V TWO GIANT ATOMS IN SEPARATE AND NESTED CONFIGURATIONS
In particular, the value of concurrence $C_{s}$ can exceed 0.5 when $\theta\rightarrow\pi/2$ and $3\pi/2$, and the concurrence $C_{N}$ is characterized by periodic oscillations ranging from zero to one at the values of $\theta$ where DFI manifests (i.e., $\theta=\pi/4,3\pi/4,5\pi/4,7\pi/4$). This indicates that two nested giant atoms in the semi-infinite waveguide system can also be prepared in the maximally entangled state. Notably, in the infinite waveguide scheme [30], it is not possible to generate entangled states with concurrence exceeding 0.5 in the separate configuration as well as maximally entangled states in the nested configuration.
VI DISCUSSION AND CONCLUSION
@@
We present a discussion on the experimental feasibility of the proposed scheme. The giant atom setup has been experimentally realized recently [29, 71, 72, 73]. In particular, Ref. [29] reports a realization that two giant atoms, whose connection points are arranged in a braided manner, are coupled to an infinite microwave waveguide. Furthermore, the coupling of two small atoms to a microwave waveguide with one end terminated by a mirror has been demonstrated in experiment [74]. These results suggest that our scheme is achievable with current and near-future experimental conditions.
@@
To summarize, we have derived the quantum master equation for multiple giant atoms coupled to a semi-infinite waveguide, where each atom has multiple connection points. We specifically analyzed the behavior of two giant atoms coupled to a semi-infinite waveguide. Concretely, we considered three configurations: separate, braided, and nested. We show that two giant atoms in braided and nested configurations can interact through the semi-infinite waveguide without decohering, i.e., there is a DFI between two giant atoms. Such DFI is not possible in semi-infinite waveguide systems containing two separate giant atoms or two small atoms. This finding goes beyond the result shown in Ref. [25] for the case of an infinite waveguide, where DFI can only be realized in the braided configuration.
@@ A.1 Two braided giant atoms with two connection points each
We derived the quantum master equation for two braided giant atoms, each coupled to a one-dimensional semi-infinite waveguide at two connection points, using the $(S,L,H)$ formalism [75, 76, 77]. In this formalism, $n$ input-output ports are needed to find the triplet for a system: S (an $n\times n$ scattering matrix), $L$ (a vector consisting of $n$ matrices that describe the system-environment interaction), and the Hamiltonian $H$ of the system (including atom-atom interaction).
@@
$=\left(\mathbf{S}_{2}\mathbf{S}_{1},\mathbf{S}_{2}\mathbf{L}_{1}+\mathbf{L}_{2},H_{1}+H_{2}+\frac{1}{2i}\left[\mathbf{L}_{2}^{\dagger}\mathbf{S}_{2}\mathbf{L}_{1}-\mathbf{L}_{1}^{\dagger}\mathbf{S}_{2}^{\dagger}\mathbf{L}_{2}\right]\right)$
@@
$G_{\text{tot}}$ $=(\mathbf{S_{tot}},\mathbf{L_{tot}},H_{tot})=G_{\text{R}}\triangleleft G_{\theta}\triangleleft G_{\text{L}}$ (22)
@@
In particular, when Eq. (A.7) is transformed into the interaction picture and the assumptions $\gamma_{1}=\gamma_{2}=\gamma_{3}=\gamma_{4}=\gamma$, $\theta_{1}=\theta_{2}=\theta_{3}=\theta$ are made, Eq. (A.7) reduces to Eq. (3) of the main text, with the coefficients given in Eq. (4).
@@ A.2 Multiple giant atoms with multiple connection points each
$+\sum_{j=1}^{M-1}\sum_{k=j+1}^{M}\sum_{n=1}^{N_{j}}\sum_{m=1}^{N_{k}}\frac{\sqrt{\gamma_{jn}\gamma_{km}}}{2}\Big{\{}\sin\left[k_{0}|x_{jn}-x_{km}|\right]+\sin\left[k_{0}(x_{jn}+x_{km})\right]\Big{\}}\left(\sigma_{+}^{(k)}\sigma_{-}^{(j)}+\text{H.c.}\right).$

### v2

@@ V TWO GIANT ATOMS IN SEPARATE AND NESTED CONFIGURATIONS
In particular, the value of concurrence $C_{eg}^{S}$ can exceed 0.5 when $\theta\rightarrow\pi/2$ and $3\pi/2$, and the concurrence $C_{eg}^{N}$ is characterized by periodic oscillations ranging from zero to one at the values of $\theta$ where DFI manifests (i.e., $\theta=\pi/4,3\pi/4,5\pi/4,7\pi/4$). This indicates that two nested giant atoms in the semi-infinite waveguide system can also be prepared in the maximally entangled state. Notably, in the infinite waveguide scheme [33], as detailed in Appendix C, it is not possible to generate entangled states with concurrence exceeding 0.5 in the separate configuration, nor to obtain maximally entangled states in the nested configuration.
VI GENERALIZATION TO MULTIPLE GIANT ATOMS
Now that we have illustrated decoherence suppression and entanglement generation for two giant atoms in a semi-infinite waveguide, it is natural to generalize the discussion to multiple atoms. Here, as shown in Fig. 6, we consider a three-giant-atom setup, with a specific focus on the braided [Fig. 6(a)(a)] and nested [Fig. 6(b)(b)] configurations that have presented superior performance in Sec. III and Sec. IV.
Figure 6: (Color online) Different configurations of three giant atoms: $(a)$ braided and $(b)$ nested. The notations here have the same meanings as those in Fig. 1.
VI.1 DFI
As could be anticipated, the DFI remains feasible for multiple giant atoms in both braided and nested configurations. To demonstrate this, we follow the approach in the above section and plot the dependence of the exchange interactions and decay rates of three braided giant atoms and three nested giant atoms on $\theta$ in Figs. 7 and 8, respectively. The relevant coefficient expressions can be obtained according to Eq. (2). By an approach analogous to the ones used in Sec. III, we conclude that the dependence of exchange interaction and decay rates for three giant atoms in braided and nested configurations on $\theta$ is a $2\pi$-periodic function. As a result, without loss of generality, our discussion is confined to the range of $\theta\in[0,2\pi]$, which already contains all relevant features of the system. For three giant atoms, in order to achieve DFI, it is necessary to have
$g_{a,b},g_{b,c},g_{a,c}$ $\neq 0,$ (16)
$\Gamma_{a},\Gamma_{b},\Gamma_{c}$ $=0,$
$\Gamma_{coll,a,b},\Gamma_{coll,b,c},\Gamma_{coll,a,c}$ $=0.$
From Figs. 7 and 8, we can see that for three giant atoms in both braided and nested configurations, there are values of $\theta$ where the exchange interaction and decay rates satisfy the condition (16), with the corresponding values given by $\theta=\pi/3,5\pi/3$ for the braided configuration and $\theta=\pi/6,\pi/2,5\pi/6,7\pi/6,3\pi/2,11\pi/6$ for the nested configuration. These results indicate that the DFI can be realized when multiple giant atoms are in the braided and nested configurations.
Note that Refs. [25, 26, 27, 28, 31] also confirmed the manifestation of DFI between multiple giant atoms, which, however, can only be observed in the braided configuration.
Figure 7: (Color online) $(a)$ Exchange interaction, $(b)$ individual decay rates, and $(c)$ collective decay rates as a function of $\theta/\pi$ for three braided giant atoms.
Figure 8: (Color online) $(a)$ Exchange interaction, $(b)$ individual decay rates, and $(c)$ collective decay rates as a function of $\theta/\pi$ for three nested giant atoms.
VI.2 Multipartite entanglement generation
We now turn to the entanglement generation of three giant atoms in braided and nested configurations. We let three giant atoms start in the state $|e\rangle_{a}|g\rangle_{b}|g\rangle_{c}$. Then the general state of the system is
$|\psi(t)\rangle$ $=c_{1}(t)|e\rangle_{a}|g\rangle_{b}|g\rangle_{c}+c_{2}(t)|g\rangle_{a}|e\rangle_{b}|g\rangle_{c}$ (17)
$\quad+c_{3}(t)|g\rangle_{a}|g\rangle_{b}|e\rangle_{c}.$
The entanglement among three giant atoms can be quantified by the linear entropy $Q$ [77], which is defined as
$Q\equiv\left[\frac{1}{n}\left(\sum_{i=1}^{n}L_{i}\right)\right],$ (18)
where $n$ denotes the size of the system, $L_{i}=2\left[1-\text{Tr}\left(\rho_{i}^{2}\right)\right]$ represents the linear entropy of the $i$-th qubit. For the three-giant-atom quantum state given by Eq. (17), the linear entropy can be expressed as [77, 78]
$Q=\frac{4}{3}\sum_{i=1}^{3}\left(|c_{i}(t)|^{2}-|c_{i}(t)|^{4}\right),$ (19)
where $c_{1}(t)$, $c_{2}(t)$, and $c_{3}(t)$ can be obtained by solving the Schrödinger equation $i\frac{\partial}{\partial t}|\psi(t)\rangle={H}_{\text{eff}}^{\prime}|\psi(t)\rangle$, with ${H}_{\text{eff}}^{\prime}$ derivable via the procedure outlined in Sec. IV.1. From Eq. (19), the $Q$ value of the state $W=\frac{1}{\sqrt{3}}\left(\lvert e_{a}g_{b}g_{c}\rangle+\lvert g_{a}e_{b}g_{c}\rangle+\lvert g_{a}g_{b}e_{c}\rangle\right)$ (a tripartite maximally entangled state) is found to be $8/9$.
In Sec. IV and Sec. V, we demonstrate that the two giant atoms can be prepared in a maximally entangled state when the phase shift is tuned to the value at which DFI manifests. However, for systems with multiple giant atoms, the situation is not always the same.
In Fig. 9, we present the time evolution of the two linear entropies $Q_{B}$ and $Q_{N}$ at the phase-shift values where the DFI manifests, where $Q_{B}$ and $Q_{N}$ characterize the three-giant-atom entanglement in braided and nested configurations, respectively. Here, we present the results for $\theta=\pi/3$ in Fig. 9(a)(a) and for $\theta=\pi/2$ in Fig. 9(b)(b) as representative examples. From Fig. 9(a)(a), it can be seen that at $\theta=\pi/3$, the peak value of $Q_{B}$ can reach 8/9, i.e., three giant atoms in the braided configuration can be prepared in the tripartite maximally entangled state–$W$ state. In Fig. 9(b)(b), for the nested configuration, when $\theta=\pi/2$, we can see that while a high degree of entanglement can be generated among the three giant atoms, the peak value of $Q_{N}$ fails to reach 8/9 required for a $W$ state. That is, in the semi-infinite waveguide system with multiple giant atoms, maximally entangled states cannot always be generated even when the DFI exists between the atoms.
Figure 9: (Color online) Linear entropies $(a)$ $Q_{B}(t)$ and $(b)$ $Q_{N}(t)$ as a function of $\gamma t$ at a given value of phase shift.
Figure 10: (Color online) Concurrences $(a)$ $C_{ee}^{B}$, $(b)$ $C_{ee}^{S}$, and $(c)$ $C_{ee}^{N}$ as a function of $\theta/\pi$ and $\gamma t$.
VII DISCUSSION
In the preceding sections, we restricted the discussion of giant-atom entanglement generation to the single-excitation subspace. We now briefly address entanglement generation for giant atoms in the two-excitation subspace. For the sake of simplicity, we focus on the case of two giant atoms. The entanglement dynamics of the two excited giant atoms across three different configurations can be obtained by numerically solving the quantum master equation (3). As shown in Fig. 10, we present the concurrences of two giant atoms with three different configurations as a function of $\theta/\pi$ and $\gamma t$ in the state $|e\rangle_{a}|e\rangle_{b}$. Figures 10(a)(a), 10(b)(b), and 10(c)(c) correspond to the braided, separate, and nested configurations, respectively. The results indicate that, compared with the infinite waveguide scheme [33], two excited giant atoms in the semi-infinite waveguide system can achieve stronger entanglement. To facilitate comparison, we also replot the entanglement dynamics of two excited giant atoms in three configurations reported in Ref. [33], shown in Appendix C. From Fig. 10 we can see that the evolution of concurrences $C_{ee}^{B}$, $C_{ee}^{S}$, and $C_{ee}^{N}$ is also periodic with a period of $2\pi$. Within one period, the maximum values of the concurrences $C_{ee}^{B}$, $C_{ee}^{S}$, and $C_{ee}^{N}$ are obtained to be $\approx$ 0.30, 0.38, and 0.38, respectively. In the infinite-waveguide system [33], as detailed in Appendix C, the maximum values of the concurrences for the two giant atoms in these three configurations are $\approx$ 0.029, 0.029, and 0.37, respectively. Namely, for the braided and separate configurations, the entanglement maximum value of two excited giant atoms in the semi-infinite waveguide system is enhanced by one order of magnitude compared with that in an infinite waveguide system.
@@ VII DISCUSSION
We next present a discussion on the experimental feasibility of the proposed scheme. The giant atom setup has been experimentally realized recently [29, 79, 80, 81]. In particular, Ref. [29] reports a realization that two giant atoms, whose connection points are arranged in a braided manner, are coupled to an infinite microwave waveguide. Furthermore, the coupling of two small atoms to a microwave waveguide with one end terminated by a mirror has been demonstrated in experiment [82]. These results suggest that our scheme is achievable with current and near-future experimental conditions.
In the single-excitation initial state, we analyzed the entanglement dynamics of two giant atoms at certain phase shift values. These phase shifts are experimentally accessible by tuning the atomic frequency, as the phase shift $\theta=k_{0}d=\omega_{0}d/v_{g}$ ($v_{g}$ represents the group velocity of the photon wave packet, $d$ is the distance between two connection points, and $\omega_{0}=\omega_{a}=\omega_{b}$ is the giant atom frequency). For example, in Ref. [29], a phase shift of $\theta=\pi/2$ is achieved when the atomic frequency is tuned to $4.645\,\text{GHz}$. Using the realistic experimental parameter $\theta=\pi/2$ reported in Ref. [29] for two braided giant atoms, the result shown in Fig. 3(b)(b) can be realized. Though Fig. 3(b)(b) is plotted up to $\gamma t=6$, the entanglement between the two braided giant atoms will keep evolving periodically between 0 and 1 when extending the time—this indicates that the entanglement can survive persistently in such periodic oscillations.
Finally, for entanglement between two atoms, it can be witnessed experimentally via violation tests of Bell’s inequalities. Considering the generalized form of Bell’s inequality [83, 84]:
$|S|\leq 2,$ (20)
where
$S=E(\vec{x},\vec{y})-E(\vec{x},\vec{y}^{\prime})+E(\vec{x}^{\prime},\vec{y})+E(\vec{x}^{\prime},\vec{y}^{\prime}),$ (21)
involves four measurements in four various orientations, and $E(\vec{x},\vec{y})$ denotes the correlation function when the two atoms are measured along orientations $\vec{x}$ and $\vec{y}$, respectively. If experimentally measured $|S|>2$, it demonstrates that non-local correlations exist between the two atoms, which indicates the presence of entanglement.
VIII CONCLUSION
@@
To summarize, we have derived the quantum master equation for multiple giant atoms coupled to a 1D semi-infinite waveguide, where each atom has multiple connection points. We specifically analyzed the behavior of two giant atoms coupled to a semi-infinite waveguide. Concretely, we considered three configurations: separate, braided, and nested. We show that two giant atoms in braided and nested configurations can interact through the semi-infinite waveguide without decohering, i.e., there is a DFI between two giant atoms. Such DFI is not possible in semi-infinite waveguide systems containing two separate giant atoms or two small atoms. This finding goes beyond the result shown in Ref. [25, 26, 27, 28, 29, 30, 31] for the case of other 1D waveguides, where DFI can only be realized in the braided configuration.
@@
Finally, we extend the discussion on DFI and entanglement generation to the system composed of multiple giant atoms coupled into a semi-infinite waveguide. We find that for multiple giant atoms in both braided and nested configurations, DFI can still be realized, and a high degree of entanglement between them can be generated.
@@ A.1 Two braided giant atoms with two connection points each
We derived the quantum master equation for two braided giant atoms, each coupled to a 1D semi-infinite waveguide at two connection points, using the $(S,L,H)$ formalism [85, 86, 87]. In this formalism, $n$ input-output ports are needed to find the triplet for a system: S (an $n\times n$ scattering matrix), $L$ (a vector consisting of $n$ matrices that describe the system-environment interaction), and the Hamiltonian $H$ of the system (including atom-atom interaction).
@@
$=\left(\mathbf{S}_{2}\mathbf{S}_{1},\mathbf{S}_{2}\mathbf{L}_{1}+\mathbf{L}_{2},H_{1}+H_{2}+\frac{1}{2i}\left[\mathbf{L}_{2}^{\dagger}\mathbf{S}_{2}\mathbf{L}_{1}-\mathbf{L}_{1}^{\dagger}\mathbf{S}_{2}^{\dagger}\mathbf{L}_{2}\right]\right).$
@@
$G_{\text{tot}}$ $=(\mathbf{S_{tot}},\mathbf{L_{tot}},H_{tot})=G_{\text{R}}\triangleleft G_{\theta}\triangleleft G_{\text{L}},$ (A.6)
@@
In particular, when Eq. (A.9) is transformed into the interaction picture and the assumptions $\gamma_{1}=\gamma_{2}=\gamma_{3}=\gamma_{4}=\gamma$, $\theta_{1}=\theta_{2}=\theta_{3}=\theta$ are made, it reduces to Eq. (3) of the main text, with the coefficients given in Eq. (4).
@@ A.2 Multiple giant atoms with multiple connection points each
$+\sum_{j=1}^{M-1}\sum_{k=j+1}^{M}\sum_{n=1}^{N_{j}}\sum_{m=1}^{N_{k}}\frac{\sqrt{\gamma_{jn}\gamma_{km}}}{2}\Big\{\sin\left[k_{0}|x_{jn}-x_{km}|\right]+\sin\left[k_{0}(x_{jn}+x_{km})\right]\Big\}\left(\sigma_{+}^{(k)}\sigma_{-}^{(j)}+\text{H.c.}\right),$
@@
Appendix C THE ENTANGLEMENT GENERATION BETWEEN TWO GIANT ATOMS WITH THREE DIFFERENT CONFIGURATIONS IN INFINITE WAVEGUIDE SYSTEMS
In this appendix, we replot the corresponding results for the infinite waveguide case, as shown in Fig. 12.
Figure 12: (Color online) Concurrences as a function of $\theta/\pi$ and $\gamma t$. In the upper and lower columns, atomic initial state $|\psi(0)\rangle=|e\rangle_{a}|g\rangle_{b}$ and $|e\rangle_{a}|e\rangle_{b}$, respectively. The two giant atoms are arranged in braided [$(a)$, $(d)$], separate [$(b)$, $(e)$], and nested configurations [$(c)$, $(f)$].
Appendix D COEFFICIENT EXPRESSIONS FOR TWO GIANT ATOMS IN SEPARATE AND NESTED CONFIGURATIONS
Table 1: Lamb shifts ($\delta\omega_{a}$, $\delta\omega_{b}$), exchange interaction ($g_{a,b}$), individual decay rates ($\Gamma_{a}$, $\Gamma_{b}$), and collective decay rates ($\Gamma_{coll,a,b}$) for two giant atoms in separate and nested configurations.
Coefficient Configuration Atom $a$ Atom $b$
$\delta\omega_{a},\delta\omega_{b}$ Separate $\gamma\sin\theta+\gamma\sin(2\theta)+\frac{\gamma}{2}\sin\theta+\frac{\gamma}{2}\sin(3\theta)$ $\gamma\sin\theta+\gamma\sin(6\theta)+\frac{\gamma}{2}\sin(5\theta)+\frac{\gamma}{2}\sin(7\theta)$
Nested $\gamma\sin(3\theta)+\gamma\sin(4\theta)+\frac{\gamma}{2}\sin\theta+\frac{\gamma}{2}\sin(7\theta)$ $\gamma\sin\theta+\gamma\sin(4\theta)+\frac{\gamma}{2}\sin(3\theta)+\frac{\gamma}{2}\sin(5\theta)$
$\Gamma_{a},\Gamma_{b}$ Separate $2\gamma+2\gamma\cos\theta+2\gamma\cos(2\theta)+\gamma\cos\theta$ $2\gamma+2\gamma\cos\theta+2\gamma\cos(6\theta)+\gamma\cos(5\theta)$
$\gamma\cos(3\theta) +$$\gamma\cos(7\theta)$
Nested $2\gamma+2\gamma\cos(3\theta)+2\gamma\cos(4\theta)+\gamma\cos\theta$ $2\gamma+2\gamma\cos\theta+2\gamma\cos(4\theta)+\gamma\cos(3\theta)$
$\gamma\cos(7\theta)$ $\gamma\cos(5\theta)$
$\Gamma_{coll,a,b}$ Separate $\gamma\!\left[\cos\theta+2\cos(2\theta)+2\cos(3\theta)+2\cos(4\theta)+\cos(5\theta)\right]$
Nested $\gamma\!\left[2\cos\theta+3\cos(2\theta)+\cos(3\theta)+\cos(5\theta)+\cos(6\theta)\right]$
$g_{a,b}$ Separate $\frac{\gamma}{2}\!\left[\sin\theta+2\sin(2\theta)+2\sin(3\theta)+2\sin(4\theta)+\sin(5\theta)\right]$
Nested $\frac{\gamma}{2}\!\left[2\sin\theta+3\sin(2\theta)+\sin(3\theta)+\sin(5\theta)+\sin(6\theta)\right]$